The Miller Corp. will pay an annual dividend of $1.50 one year from now.

  • Analysts expect this dividend to grow at 12% per year thereafter until the 3rd year.
  • After then, growth will level off at 4% per year forever.

What is the value of its stock today if the firm’s required return is 20%?
Answer #1

D1 =    1.5      
          
D2 =1.5*(1+12%)   1.68      
D3 = 1.68*(1+12%)=   1.8816      
D4 = 1.8816*(1+4%)=   1.956864      
          
Price of stock at end of Year 3 (P3) = D4/(Ke -g)          
1.956864   /(20%-4%)      
12.2304          
          
Current market price today is present value of stock that is equal to present value of dividend received upto 2 Years and Price of stock received at end of 2 Year          
          
Year   Cash flows   P.V.F.@ 20%   Cash flows * PVF
1 (D1)   1.5   0.8333333333   1.25
2 (D2)   1.68   0.6944444444   1.166666667
3(D3)   1.8816   0.5787037037   1.088888889
3 (P3)   12.2304   0.5787037037   7.077777778
          
Current value 10.58333333

So Current price of stock is $10.5833